# COOL: Efficient and Reliable Chain-Oriented Objective Logic with Neural Networks Feedback Control

### v3

@@ Appendix F Implementation Toolchain
To fully implement the CoL DSL and adapt it to NNFC, we choose to build COOL from the ground up rather than extending existing DSL frameworks such as Xtext Bettini (2016) or Groovy King (2020). We use C++ as the primary language to meet the execution efficiency requirements for the numerous tree operations inherent in the DSL program synthesis process. For development efficiency, we utilize Lex Lesk & Schmidt (1975) and YACC Johnson et al. (1975) for syntax and semantic parsing, respectively. The neural network components are implemented in Python, leveraging the PyTorch library Imambi et al. (2021) to support machine learning tasks effectively. Table 8 shows the detailed code effort involved in developing the different components of COOL across various programming languages.
@@ Appendix G Optimization strategy
    Gradient-Based Regulation: We employ gradient-based regulation, a widely used strategy in program synthesis Cui & Zhu (2021); Liang et al. (2018); Chaudhuri et al. (2021). By evaluating the slope or rate of change between consecutive stages, gradients help us make dynamic adjustments to synthesis paths. In our approach, we regulate skipping by applying a gradient to the heuristic values at each stage in the CoL. We encourage skipping when the heuristic gradient from one stage to the next is positive. Conversely, if the gradient is negative, we suppress skipping.
@@ Appendix H Future work
In future work, we aim to enhance the capability of the COOL framework by exploring the implementation of CoL and NNFC in more complex scenarios, such as managing dependencies among DSL libraries and object-oriented development. We plan to facilitate community collaboration by developing more DSL libraries to expand COOL’s applications. Additionally, we are interested in integrating COOL with language models. As these models evolve, ensuring ethical and accurate reasoning becomes increasingly crucial Jacovi & Goldberg (2020); Chen et al. (2022); Li et al. (2022). The COOL framework, including CoL’s constraints on rule application and NNFC’s structured agent interactions, helps to enhance reasoning faithfulness, preventing harmful reasoning logic. We hope our work will serve as a reliable bridge for interaction and understanding between human cognitive processes and language model reasoning.
@@ Appendix I CoL DSL for Relational Tasks (Source Code)
Neural Search Optimization: Neural networks are key for optimizing search in program synthesis. Projects like Kalyan et al. (2018); Zhang et al. (2023) and Li et al. (2024) use neural networks to provide oracle-like guidance, while Neo Feng et al. (2018), Flashmeta Polozov & Gulwani (2015), and Concord Chen et al. (2020) prune search spaces with infeasible partial programs. COOL employs both strategies to enhance efficiency.
Multi-step Program Synthesis: Chain-of-Thought (CoT) Wei et al. (2022) enhances LLMs by breaking tasks into subtasks. Projects like Zhou et al. (2022); Shi et al. (2023b) and Zheng et al. (2023) use this in program synthesis. Compared to CoT, which directly decomposes tasks, CoL does so indirectly by constraining rule applications.
Reinforcement Learning: Reinforcement learning improves neural agents in program synthesis through feedback, as seen in Eberhardinger et al. (2023); Liu et al. (2024); Bunel et al. (2018), Concord Chen et al. (2020), and Quiet-STaR Zelikman et al. (2024). NNFC similarly refines control flow but serves an auxiliary role for programmer strategies in synthesis rather than dominating it.

### v4

@@ Appendix F Implementation Toolchain
To fully implement the CoL DSL and adapt it to NNFC, we choose to build COOL from the ground up rather than extending existing DSL frameworks such as Xtext (Bettini, 2016) or Groovy (King, 2020). We use C++ as the primary language to meet the execution efficiency requirements for the numerous tree operations inherent in the DSL program synthesis process. For development efficiency, we utilize Lex (Lesk & Schmidt, 1975) and YACC (Johnson et al., 1975) for syntax and semantic parsing, respectively. The neural network components are implemented in Python, leveraging the PyTorch library (Imambi et al., 2021) to support machine learning tasks effectively. Table 8 shows the detailed code effort involved in developing the different components of COOL across various programming languages.
@@ Appendix G Optimization strategy
    Gradient-Based Regulation: We employ gradient-based regulation, a widely used strategy in program synthesis (Cui & Zhu, 2021; Liang et al., 2018; Chaudhuri et al., 2021). By evaluating the slope or rate of change between consecutive stages, gradients help us make dynamic adjustments to synthesis paths. In our approach, we regulate skipping by applying a gradient to the heuristic values at each stage in the CoL. We encourage skipping when the heuristic gradient from one stage to the next is positive. Conversely, if the gradient is negative, we suppress skipping.
@@ Appendix H Future Work
In future work, we aim to enhance the capability of the COOL framework by exploring the implementation of CoL and NNFC in more complex scenarios, such as managing dependencies among DSL libraries and object-oriented development. We plan to facilitate community collaboration by developing more DSL libraries to expand COOL’s applications. Additionally, we are interested in integrating COOL with language models. As these models evolve, ensuring ethical and accurate reasoning becomes increasingly crucial (Jacovi & Goldberg, 2020; Chen et al., 2022; Li et al., 2022). The COOL framework, including CoL’s constraints on rule application and NNFC’s structured agent interactions, helps to enhance reasoning faithfulness, preventing harmful reasoning logic. We hope our work will serve as a reliable bridge for interaction and understanding between human cognitive processes and language model reasoning.
@@ Appendix I CoL DSL for Relational Tasks (Source Code)
Neural Search Optimization: Neural networks are key for optimizing search in program synthesis. Projects like (Kalyan et al., 2018; Zhang et al., 2023) and (Li et al., 2024) use neural networks to provide oracle-like guidance, while Neo (Feng et al., 2018), Flashmeta (Polozov & Gulwani, 2015), and Concord (Chen et al., 2020) prune search spaces with infeasible partial programs. COOL employs both strategies to enhance efficiency.
Multi-step Program Synthesis: Chain-of-Thought (CoT) (Wei et al., 2022) enhances LLMs by breaking tasks into subtasks. Projects like (Zhou et al., 2022; Shi et al., 2023b) and (Zheng et al., 2023) use this in program synthesis. Compared to CoT, which directly decomposes tasks, CoL does so indirectly by constraining rule applications.
Reinforcement Learning: Reinforcement learning improves neural agents in program synthesis through feedback, as seen in (Eberhardinger et al., 2023; Liu et al., 2024; Bunel et al., 2018), Concord (Chen et al., 2020), and Quiet-STaR (Zelikman et al., 2024). NNFC similarly refines control flow but serves an auxiliary role for programmer strategies in synthesis rather than dominating it.